1.Shen took out a loan for 292 days and was charged simple interest at an annual rate of 2.5%. The total interest he paid on the loan was $146. How much money did Shen borrow? Assume that there are 365 days in a year, and do not round any intermediate computations.

2. To purchase $12,600 worth of lab equipment for her business, Isabel made a down payment of

$1900 and took out a business loan for the rest. After 2 years of paying monthly payments of

$474.24, she finally paid off the loan.

a) What was the total amount Isabel ended up paying for the equipment (including the down payment and monthly payments)?

b) How much interest did Isabel pay on the loan?
Answer #1

1:

Interest = Principal*Number of periods*Rate per period

146= Principal*292* 0.025 /365

146= Principal *0.02

Principal = $7300

2:

Loan amount = Cost of equipment- down payment

= 12600-1900

=10700

A: Total amount paid = Instalment amount*Number

= 474.24*2*12

=11381.76

B: Interest on the loan = Total amount paid- Loan taken

=11381.76- 10700

=681.76